The magnetic field energy in an inductor changes from maximum to minimum value in 5.0ms when connected to AC source. Find the frequency of the source

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
=>min=0,max=peak
=>4*5ms=f
=>f=50Hz


Arun Kumar
IIT Delhi
Askiitians Faculty
Chandra Sekhara Rao
7 years ago
very good answer is followed by time period. for a quater part of wave pulse energy going down to maximum value to minimum vale. i,e
T/4 = 5ms
T =20ms
f = 1/T  =1/20ms = 50Hz